we have a cart with a total mass of 1014kg going down a ramp of 7.6 degrees with a net force on the system pf 1.28N. what would be the F theoretical(N) and the Y theoretical (N)?
Answer #1

a)

Fx = F cos x = 1.28* cos 7.6 =1.269 N

B)

Fy = N = mg cos x = 1014 * 9.8 cos 7.6 = 9850 N

======

Comment before rate in case any doubt, will reply for sure.. goodluck